Can someone help me with this problem?Solve the following initial value problem de 23 (2)+3 (do (z)) – 10y (2) = -24 el–22) with y(0) = 2, dy(0) dc = -18
Answer #1

--- da2 with d2 yca) + 3 Cola y («)) -10 y) = 1-24 0-24 yco)=2, dyco) -18 da ted y=ema be a treicel solution, then aunitary eIA = 14 Dang ♡ x2 + 2A + 2B=0 5A-2B = 14 + DA= 2 From ☺ B =-A = -2 Hence the required solution of or YCH) - 20-59_2222 +2e-2X